OpenStudy (macaylam):

Mira is playing a game where she is trying to toss a ball into a large basket or a small basket. The large basket is worth 2 points, and the small basket is worth 3 points. Mira made 15 baskets for a total of 35 points. If x represents the number of large baskets made and y represents the number of small baskets made, how many large baskets did she make? x+y=15 2x+3y=35 2 baskets 5 baskets 7 baskets 10 baskets I think its 10

OpenStudy (cp123):

x=15-y. Substitute that into your second equation 2(15-y)+3y=35. Simplify. 30-2y+3y=35. combine like terms. 30-y=35. subtract. 30 from both sides, then you get y=5

OpenStudy (cp123):

Now substitute that into both equations and solve for x. x+5=15. subtract 5 from both sides. you get x=10. Now put both 5 and 10 into the second equation to see if it is true.

OpenStudy (cp123):

10+5=15 true. 2(10)+3(5)=20+15=35. true! so x=10 and y=5

OpenStudy (cp123):

also i meant to do +y when combining like terms, not -y. sorry about that

OpenStudy (macaylam):

thanks so much:)